Three children are sitting on a seasaw in such a way that it balances. A 20kg and 30kg boy are on opposite sides at a distance B and 3B , respectively from the pivot. The rod is held horizontal and then released. Find its angular acceleration at the instant it is released.
1
Expert's answer
2020-05-04T12:43:59-0400
"\tau=30g(3b)-20gb=70gb"

"I=30(3b)^2-20b^2=250b^2"

"\alpha=\frac{\tau}{I}=\frac{70gb}{250b^2}=\frac{70(10)}{250}=2.8\frac{rad}{s^2}"